OpenStudy (anonymous):

I'm supposed to Find the derivative of f(x) = 12x2 + 8x at x = 9.

OpenStudy (anonymous):

I've seen the problem done before, but I'm just confused on the process Someone said f(x) = 12x^2 + 8x f(9) = 12(9) + 8 f(9) = 224 but i don't understand what they did with the extra 'x'

OpenStudy (jdoe0001):

I gather they meant the "slope" at x = 9 in which case, it means, use the "power rule" to get the derivative of that then in the derivative function, set x = 9

OpenStudy (anonymous):

I'm not sure I follow. the derivative function is just (the original one)/x ?

OpenStudy (jdoe0001):

are you doing calc?

OpenStudy (anonymous):

Precalc, sir/ma'am

OpenStudy (jdoe0001):

hmmm.... oddly enough.... usually that'd be in calc well, http://www.math-prof.com/Calculus_1/Calc_Ch_06.asp <--- the power rule is atop

OpenStudy (jdoe0001):

derivative of a constant is 0

OpenStudy (anonymous):

Alright, so what I'm seeing is the x^2 (x^n-1) become x^1, and turn x^2 into 2x^1?

OpenStudy (jdoe0001):

\(\bf f(x) = 12x^2 + 8x \implies \cfrac{dy}{dx} = (12 \times 2) x^{2-1} + (8 \times 1) x^{1-1}\\\quad \\ \cfrac{dy}{dx} = 24x+8\\\quad \\ \textit{setting } x =5 \textit{ will give the slope at that point}\)

OpenStudy (anonymous):

Wonderful! Thank you!

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

How might this rule apply with 'x' as a denominator? I suspect it would remain the same since there is no power value?